# A Robust Ensemble Algorithm for Ischemic Stroke Lesion Segmentation: Generalizability and Clinical Utility Beyond the ISLES Challenge

### v1

@@ Introduction
Brain imaging is crucial to evaluate tissue viability and fate in ischemic stroke. Magnetic resonance imaging (MRI) supports physicians across diverse stages of the disease. It helps define the optimal reperfusion treatment, unveils the stroke etiology, and sheds light on prognostic clinical outcomes [czap2021overview]. Diffusion-weighted imaging (DWI) is considered the current gold standard for imaging the ischemic core [goyal2020challenging, giancardo2023segmentation]. Although imperfect, DWI is the only imaging technique reliably demonstrating parenchymal injury within minutes to hours from the stroke onset [merino2010imaging]. Currently, deep learning algorithms are revolutionizing medical imaging, demonstrating unprecedented performance across multiple radiological tasks. Segmentation of ischemic stroke tissue using deep learning has been proposed in different works [chen2017fully, federau2020improved, liu2021deep, alis2021inter, nazari2023predicting]. The complexity of the task lies in multiple sources of variability that involve image- (e.g., driven by center- and scanner-specific MRI acquisition differences, artifacts mimicking ischemic lesions [roberts2003diffusion], time-dependent DWI signaling [merino2010imaging, chalela2007magnetic], etc.), patient- (e.g., age [fung2011mr]) and disease-specific characteristics (such as the subtype of stroke and its etiology [kang2003association]). Little is known, however, about the real-world transferability potential of deep learning algorithms for ischemic stroke segmentation, their generalization towards diverse cohorts and image characteristics, and their ultimate clinical utility.
@@
Biomedical challenges are international competitions aiming to benchmark task-specific algorithms under controlled settings [maier2018rankings]. The organization of medical image challenges has rapidly grown, enabling to tackle problems related to diverse organs, tasks (e.g., lesion detection or anatomy segmentation), and image modalities (such as MRI, CT, among others) [menze2014multimodal, litjens2014evaluation, sekuboyina2021verse, sirinukunwattana2017gland, bilic2023liver]. Challenges are now considered a de facto gold standard for algorithm comparison by the research community [antonelli2022medical] and have also been adopted by the Radiological Society of North America (RSNA)aaahttps://www.rsna.org/rsnai/ai-image-challenge. Segmentation of stroke lesions from MRI has not been an exception, and the number of methods devised targeting this task considerably increased following the 2015 Ischemic Stroke Lesion Segmentation (ISLES) challenge [maier2017isles]. ISLES’15 is considered a reference evaluation tool for the segmentation of brain ischemia. In the past few years, studies highlighting the strengths and weaknesses of challenge organization emerged, providing good implementation practices [reinke2018exploit, maier2018rankings, maier2020bias]. Such initiatives considerably improved the quality of current challenges in terms of execution, interpretation, fairness, transparency, and reproducibility.
@@
Biomedical challenges, when properly designed, are powerful. They operate as international problem-solving sprints that involve leading researchers worldwide. Therefore, we take advantage of such an event to rapidly prototype and identify candidate ischemic stroke segmentation algorithms. We hypothesized that (1) a challenge might yield an algorithm or a strategy that reliably detects and segments brain ischemia under real-world, heterogeneous data scenarios, and (2) such an algorithm may generalize above the challenge context to real-world data, thus becoming relevant to downstream clinical analysis. We organized the ISLES’22 challenge during the 2022 International Conference on Medical Image Computing and Computer-Assisted Intervention (MICCAI) [GC_isles, islesdocument] to test these hypotheses. ISLES’22 builds on top of the experience gained from the earlier ISLES’15, overcoming some of its drawbacks by adhering to current challenge standards [reinke2018exploit, maier2020bias], by using a standardized platform [GC_isles] for the fair assessment of software solutions, and by including more than six times the number of patients than ISLES’15. This study describes the organization of the challenge, the development of a robust deep learning ensemble solution capable of generalizing to multiple axes of data variability, its validation using the largest available (external) stroke dataset, and, ultimately, an illustration of its practical utility in facilitating downstream clinical analysis on real-world data. The contribution of this study is two-fold. First, we introduce an algorithm capable of identifying and segmenting ischemic stroke in MRI at a level comparable to expert radiologists. Furthermore, through a Turing-like experiment, we found that the algorithm produces superior qualitative segmentations, which neuroradiologists prefer over manually delineated expert counterparts. Secondly, our investigation is pioneering in reporting a challenge-derived algorithm whose (clinical) utility and applicability transcend the primary challenge objectives, showcasing its usefulness for subsequent clinical analyses.
@@ Results
Figure 1(A) summarizes the ISLES’22 challenge. In the model development phase, teams utilized labeled datasets to create algorithmic solutions, which were evaluated using undisclosed data in the model testing phase. Differently from other imaging challenges, the challenge datasets were kept raw and un-processed to emulate a real-world setting, thus challenging participants to devise end-to-end algorithmic pipelines. Figure 1(A) also showcases an overview of the external dataset used to evaluate algorithmic performance in a post-challenge, ‘real-world’ data setting. The ISLES’22 challenge was held during May-August 2022. By the challenge closing date, there were 476 registered participants and 325 dataset downloads. A total of 20 teams tested their algorithms on the remote server using the sanity-check phase. Finally, we received 15 submissions to the challenge’s final test-phase, from which 12 fulfilled the ISLES’22-MICCAI requirements [islesdocument]. The 12 participating teams submitted deep learning-based solutions. In Figure 1(B), a fingerprint of the methods is shown. nnUnet [isensee2021nnu] and UNet-like [ronneberger2015u, cciccek20163d] neural networks were the most prevalent architectures. The most frequently used loss function was a combination of Dice with categorical cross-entropy. Figure 1(C) shows a pictorial ranking of the algorithmic submissions. Most teams preferred to use all the available MRI modalities (DWI, apparent diffusion coefficient ‘ADC’, and fluid-attenuated inversion recovery ‘FLAIR’) as inputs to the model. It is worth observing that the top-3 ranked teams used diverse deep learning pipelines in terms of architecture, loss function, and model inputs.
@@
Team ‘SEALS’. The participants utilized DWI and ADC images as input for their algorithm. Image pre-processing involved the resampling of the scans to a 1$\times$1$\times$1 mm3 voxel resolution, followed by a z-score image normalization. The nn-UNet pipeline [isensee2021nnu] was employed for training a 3D full-resolution UNet. A 1/7 subset of the training dataset was held out to evaluate the performance of the model. The remaining 6/7 parts of the dataset were used to train models through 5-fold cross-validation. A combined Dice loss with categorical cross-entropy was used. Data augmentation transforms were used, including image flipping and Gaussian noise addition. The final submission to the challenge was an ensemble of the five trained models. The algorithm is publicly available at https://github.com/Tabrisrei/ISLES22_SEALS.
@@
Team ‘NVAUTO’. The team proposed an automated 3D semantic segmentation solution implemented with Auto3DSeg [auto3dseg]. The algorithm automated most deep learning steps and decision choices. DWI and ADC images were used as input to the model after voxel resampling to 1$\times$1$\times$1 mm3 and z-score normalization. SegResNet [myronenko20193d] models were trained through 5-fold cross-validation. Several augmentation transforms were used, including flipping, rotation, scaling, smoothing, intensity-scaling and -shifting. Random cropped patches of dimensions 192$\times$192$\times$128 were adopted. The models were trained on an 8-GPU NVIDIA V100 machine with an AdamW optimizer and unitary batch size. Dice loss with focal loss using deep supervision was used as a loss function. The model was first pre-trained on the BRATS 2021 dataset [baid2021rsna]. The final algorithm was an ensemble of 15 models obtained through a 3-time 5-fold cross-validation. The algorithm is publicly available at https://github.com/mahfuzmohammad/isles22.
@@
Team ‘SWAN’. The participants used the Factorizer [ashtari2023factorizer] algorithm to construct an end-to-end, linearly scalable model for stroke lesion segmentation. Factorizer is a family of models that leverage the power of Non-negative Matrix Factorization (NMF) to extract discriminative and meaningful feature maps. The algorithm uses a differentiable NMF layer that can be back-propagated during the training of deep learning models. A Factorizer block is constructed by replacing the self-attention layer of a vision transformer block [dosovitskiy2020image] with an NMF-based module and then integrating them into a U-shaped architecture with skip connections. The participants used a Swin Factorizer, which combines NMF with the shifted-window idea inspired by Liu et al. [liu2021swin] to effectively exploit local context. Preprocessing involves FLAIR-to-DWI image registration using Elastix [klein2009elastix] and z-score normalization. Various data augmentation techniques were performed, including random affine transforms, flips, and random intensity scalings. Deep supervision was used at the three highest decoder resolution levels for training the models. The final challenge submission was an ensemble of Swin Factorizers and UNet models with residual blocks [he2016deep] (a.k.a ResU-Net) obtained through 5-fold cross-validation. The pipeline and pre-trained models are available at https://github.com/pashtari/factorizer-isles22.
@@
Algorithmic robustness to out-of-domain data (unseen during the model’s development phase) is crucial for evaluating the algorithm’s transferability to real-world centers. Figure 3(A) shows how the ensemble model performs over test-phase data from seen (centers ##\##1 and ##\##2) and unseen (center ##\##3) centers during the development of the algorithm. The distribution of the metrics obtained over the unseen center ##\##3 is similar to the metric’s distribution obtained over the seen center ##\##1, suggesting an overall good generalization to new center data (Dice p-value = 0.73, F1 score p-value = 0.60, ALD p-value = 0.42, AVD p-value = 0.08, Wilcoxon rank-sum tests). The performance obtained over the seen center ##\##2 is lower in terms of Dice score compared to the center ##\##1 (p-value < 0.001, Wilcoxon rank-sum test). The F1 score, AVD and ALD metrics are similar between centers ##\##1 and ##\##2 (F1 score p-value = 0.60, ALD p-value = 0.26, AVD p-value = 0.28, Wilcoxon rank-sum tests). The lower Dice scores in center ##\##2 can be explained by two cohort confounders. Firstly, the scans from center ##\##2 include smaller lesion volumes than scans from the other two centers [hernandez2022isles] (p-value = 0.039 for center #$\#1$ vs center #$\#2$, p-value = 0.001 for center #$\#2$ vs center #$\#3$, p-value = 0.56 for center #$\#1$ vs center #$\#3$, Wilcoxon rank-sum tests). Figure S5 (see Supplementary Materials) shows the non-linear, monotonic correlation between lesion size and Dice scores for the test set data. The fact that larger objects (i.e., brain lesions) benefit from higher Dice scores is well known and, therefore, is associated with the found results [maier2022metrics, reinke2023understanding]. Secondly, unlike the train phase data, which considers scans acquired in the sub-acute stroke phase after reperfusion treatment, the test set scans from center ##\##2 are acquired in the acute stroke phase, before the patient’s reperfusion treatment, which is known to be a harder task for the algorithms [maier2017isles]. The following sections analyze both of the aforementioned confounding factors. It is worth noting a potential third confounder related to the imbalance in training data (4:1 for centers #1:#2, as shown in Table 4). However, we chose to disregard this factor because the model demonstrated strong generalization abilities to scans from an entirely new center.
@@
Two expert neuroradiologists annotated ten randomly sampled scans from the ISLES’22 training set [hernandez2022isles]. When comparing their delineations against the ground truth masks, they achieved a median $\pm$ interquartile range Dice score of 0.92 $\pm$ 0.16 and a lesion-wise F1 score of 0.82 $\pm$ 0.30. The ensemble algorithm yielded, over the entire test set, a Dice score of 0.82 $\pm$ 0.12 and an F1 score of 0.86 $\pm$ 0.21. Figure 5 shows the median Dice MRI scan (or the closest to the median) for diverse stroke scenarios. The visual results of the predicted ischemic masks suggest that the ensemble predictions (green delineations) closely follow the manually segmented lesions (red delineations), exposing the algorithm’s capability to cope with diverse types of stroke patterns and configurations. The quantitative and qualitative results suggest robustness towards heterogeneous clinical and imaging scenarios.
@@
In a Turing-like test, nine experienced radiologists rated the stroke segmentation quality of the ISLES’22 test data. Radiologists received forty or forty-one randomized images with each image being delineated either by an expert or by the ensemble algorithm and were asked to rate the completeness and correctness of the lesion masks in a 1-to-6 (worst-to-best) scale. Boxplots of this experiment are shown in Fig. 6. Interestingly, the ensemble algorithm exhibits statistically significantly higher ratings than the experts (p-value = 0.02 when considering the segmentation completeness and p-value < 0.001 when considering the segmentation correctness, Wilcoxon signed-rank tests). The observation that experts find deep learning segmentations to be qualitatively superior to manually traced ones is unsurprising, given that similar findings were reported in prior research [kofler2021we].
@@
We tested the ISLES’22 ensemble algorithm over the largest available acute and early sub-acute ischemic stroke cohort (N = 1686) retrieved from the Johns Hopkins Comprehensive Stroke Center [liu2023large]. The data were collected over ten years utilizing eleven magnetic field scanners operating at either 1.5T or 3T, sourced from four different vendors. Due to continual technological advancements over the last decade, the dataset comprises diverse MRI acquisition protocols and machine technologies, leading to strong variations in image quality (such as resolution, signal-to-noise ratio, and contrast-to-noise ratio). This makes this dataset uniquely suited to evaluate the generalizability of algorithms. Ischemic lesion masks were annotated by the dataset authors through manual delineation. Patient age was 62.5 $\pm$ 13.3 years. 907 patients (53.8%) were male. Reported race or ethnicity included 753 (44.7%) Black or African American, 490 (29.1%) White, and 40 (2.4%) Asian. 876 (52.0%) MRI scans were acquired after thrombolysis treatment with intravenous tissue-type plasminogen activator.

### v2

@@ Introduction
Brain imaging is crucial to evaluate tissue viability and fate in ischemic stroke. Magnetic resonance imaging (MRI) supports physicians across diverse stages of the disease. It helps define the optimal reperfusion treatment, unveils the stroke etiology, and sheds light on prognostic clinical outcomes [1]. Diffusion-weighted imaging (DWI) is considered the current gold standard for imaging the ischemic core [2, 3]. Although imperfect, DWI is the only imaging technique reliably demonstrating parenchymal injury within minutes to hours from the stroke onset [4]. Currently, deep learning algorithms are revolutionizing medical imaging, demonstrating unprecedented performance across multiple radiological tasks. Segmentation of ischemic stroke tissue using deep learning has been proposed in different works [5, 6, 7, 8, 9]. The complexity of the task lies in multiple sources of variability that involve image- (e.g., driven by center- and scanner-specific MRI acquisition differences, artifacts mimicking ischemic lesions [10], time-dependent DWI signaling [4, 11], etc.), patient- (e.g., age [12]) and disease-specific characteristics (such as the subtype of stroke and its etiology [13]). Little is known, however, about the real-world transferability potential of deep learning algorithms for ischemic stroke segmentation, their generalization towards diverse cohorts and image characteristics, and their ultimate clinical utility.
@@
Biomedical challenges are international competitions aiming to benchmark task-specific algorithms under controlled settings [14]. The organization of medical image challenges has rapidly grown, enabling to tackle problems related to diverse organs, tasks (e.g., lesion detection or anatomy segmentation), and image modalities (such as MRI, CT, among others) [15, 16, 17, 18, 19]. Challenges are now considered a de facto gold standard for algorithm comparison by the research community [20] and have also been adopted by the Radiological Society of North America (RSNA)aaahttps://www.rsna.org/rsnai/ai-image-challenge. Segmentation of stroke lesions from MRI has not been an exception, and the number of methods devised targeting this task considerably increased following the 2015 Ischemic Stroke Lesion Segmentation (ISLES) challenge [21]. ISLES’15 is considered a reference evaluation tool for the segmentation of brain ischemia. In the past few years, studies highlighting the strengths and weaknesses of challenge organization emerged, providing good implementation practices [22, 14, 23]. Such initiatives considerably improved the quality of current challenges in terms of execution, interpretation, fairness, transparency, and reproducibility.
@@
Biomedical challenges, when properly designed, are powerful. They operate as international problem-solving sprints that involve leading researchers worldwide. Therefore, we take advantage of such an event to rapidly prototype and identify candidate ischemic stroke segmentation algorithms. We hypothesized that (1) a challenge might yield an algorithm or a strategy that reliably detects and segments brain ischemia under real-world, heterogeneous data scenarios, and (2) such an algorithm may generalize above the challenge context to real-world data, thus becoming relevant to downstream clinical analysis. We organized the ISLES’22 challenge during the 2022 International Conference on Medical Image Computing and Computer-Assisted Intervention (MICCAI) [24, 25] to test these hypotheses. ISLES’22 builds on top of the experience gained from the earlier ISLES’15, overcoming some of its drawbacks by adhering to current challenge standards [22, 23], by using a standardized platform [24] for the fair assessment of software solutions, and by including more than six times the number of patients than ISLES’15. This study describes the organization of the challenge, the development of a robust deep learning ensemble solution capable of generalizing to multiple axes of data variability, its validation using the largest available (external) stroke dataset, and, ultimately, an illustration of its practical utility in facilitating downstream clinical analysis on real-world data. The contribution of this study is two-fold. First, we introduce an algorithm capable of identifying and segmenting ischemic stroke in MRI at a level comparable to expert radiologists. Furthermore, through a Turing-like experiment, we found that the algorithm produces superior qualitative segmentations, which neuroradiologists prefer over manually delineated expert counterparts. Secondly, our investigation is pioneering in reporting a challenge-derived algorithm whose (clinical) utility and applicability transcend the primary challenge objectives, showcasing its usefulness for subsequent clinical analyses.
@@ Results
Figure 1(A) summarizes the ISLES’22 challenge. In the model development phase, teams utilized labeled datasets to create algorithmic solutions, which were evaluated using undisclosed data in the model testing phase. Differently from other imaging challenges, the challenge datasets were kept raw and un-processed to emulate a real-world setting, thus challenging participants to devise end-to-end algorithmic pipelines. Figure 1(A) also showcases an overview of the external dataset used to evaluate algorithmic performance in a post-challenge, ‘real-world’ data setting. The ISLES’22 challenge was held during May-August 2022. By the challenge closing date, there were 476 registered participants and 325 dataset downloads. A total of 20 teams tested their algorithms on the remote server using the sanity-check phase. Finally, we received 15 submissions to the challenge’s final test-phase, from which 12 fulfilled the ISLES’22-MICCAI requirements [25]. The 12 participating teams submitted deep learning-based solutions. In Figure 1(B), a fingerprint of the methods is shown. nnUnet [26] and UNet-like [27, 28] neural networks were the most prevalent architectures. The most frequently used loss function was a combination of Dice with categorical cross-entropy. Figure 1(C) shows a pictorial ranking of the algorithmic submissions. Most teams preferred to use all the available MRI modalities (DWI, apparent diffusion coefficient ‘ADC’, and fluid-attenuated inversion recovery ‘FLAIR’) as inputs to the model. It is worth observing that the top-3 ranked teams used diverse deep learning pipelines in terms of architecture, loss function, and model inputs.
@@
Team ‘SEALS’. The participants utilized DWI and ADC images as input for their algorithm. Image pre-processing involved the resampling of the scans to a 1$\times$1$\times$1 mm3 voxel resolution, followed by a z-score image normalization. The nn-UNet pipeline [26] was employed for training a 3D full-resolution UNet. A 1/7 subset of the training dataset was held out to evaluate the performance of the model. The remaining 6/7 parts of the dataset were used to train models through 5-fold cross-validation. A combined Dice loss with categorical cross-entropy was used. Data augmentation transforms were used, including image flipping and Gaussian noise addition. The final submission to the challenge was an ensemble of the five trained models. The algorithm is publicly available at https://github.com/Tabrisrei/ISLES22_SEALS.
@@
Team ‘NVAUTO’. The team proposed an automated 3D semantic segmentation solution implemented with Auto3DSeg [29]. The algorithm automated most deep learning steps and decision choices. DWI and ADC images were used as input to the model after voxel resampling to 1$\times$1$\times$1 mm3 and z-score normalization. SegResNet [30] models were trained through 5-fold cross-validation. Several augmentation transforms were used, including flipping, rotation, scaling, smoothing, intensity-scaling and -shifting. Random cropped patches of dimensions 192$\times$192$\times$128 were adopted. The models were trained on an 8-GPU NVIDIA V100 machine with an AdamW optimizer and unitary batch size. Dice loss with focal loss using deep supervision was used as a loss function. The model was first pre-trained on the BRATS 2021 dataset [31]. The final algorithm was an ensemble of 15 models obtained through a 3-time 5-fold cross-validation. The algorithm is publicly available at https://github.com/mahfuzmohammad/isles22.
@@
Team ‘SWAN’. The participants used the Factorizer [32] algorithm to construct an end-to-end, linearly scalable model for stroke lesion segmentation. Factorizer is a family of models that leverage the power of Non-negative Matrix Factorization (NMF) to extract discriminative and meaningful feature maps. The algorithm uses a differentiable NMF layer that can be back-propagated during the training of deep learning models. A Factorizer block is constructed by replacing the self-attention layer of a vision transformer block [33] with an NMF-based module and then integrating them into a U-shaped architecture with skip connections. The participants used a Swin Factorizer, which combines NMF with the shifted-window idea inspired by Liu et al. [34] to effectively exploit local context. Preprocessing involves FLAIR-to-DWI image registration using Elastix [35] and z-score normalization. Various data augmentation techniques were performed, including random affine transforms, flips, and random intensity scalings. Deep supervision was used at the three highest decoder resolution levels for training the models. The final challenge submission was an ensemble of Swin Factorizers and UNet models with residual blocks [36] (a.k.a ResU-Net) obtained through 5-fold cross-validation. The pipeline and pre-trained models are available at https://github.com/pashtari/factorizer-isles22.
@@
Algorithmic robustness to out-of-domain data (unseen during the model’s development phase) is crucial for evaluating the algorithm’s transferability to real-world centers. Figure 3(A) shows how the ensemble model performs over test-phase data from seen (centers ##\##1 and ##\##2) and unseen (center ##\##3) centers during the development of the algorithm. The distribution of the metrics obtained over the unseen center ##\##3 is similar to the metric’s distribution obtained over the seen center ##\##1, suggesting an overall good generalization to new center data (Dice p-value = 0.73, F1 score p-value = 0.60, ALD p-value = 0.42, AVD p-value = 0.08, Wilcoxon rank-sum tests). The performance obtained over the seen center ##\##2 is lower in terms of Dice score compared to the center ##\##1 (p-value < 0.001, Wilcoxon rank-sum test). The F1 score, AVD and ALD metrics are similar between centers ##\##1 and ##\##2 (F1 score p-value = 0.60, ALD p-value = 0.26, AVD p-value = 0.28, Wilcoxon rank-sum tests). The lower Dice scores in center ##\##2 can be explained by two cohort confounders. Firstly, the scans from center ##\##2 include smaller lesion volumes than scans from the other two centers [37] (p-value = 0.039 for center #$\#1$ vs center #$\#2$, p-value = 0.001 for center #$\#2$ vs center #$\#3$, p-value = 0.56 for center #$\#1$ vs center #$\#3$, Wilcoxon rank-sum tests). Figure S5 (see Supplementary Materials) shows the non-linear, monotonic correlation between lesion size and Dice scores for the test set data. The fact that larger objects (i.e., brain lesions) benefit from higher Dice scores is well known and, therefore, is associated with the found results [38, 39]. Secondly, unlike the train phase data, which considers scans acquired in the sub-acute stroke phase after reperfusion treatment, the test set scans from center ##\##2 are acquired in the acute stroke phase, before the patient’s reperfusion treatment, which is known to be a harder task for the algorithms [21]. The following sections analyze both of the aforementioned confounding factors. It is worth noting a potential third confounder related to the imbalance in training data (4:1 for centers #1:#2, as shown in Table 4). However, we chose to disregard this factor because the model demonstrated strong generalization abilities to scans from an entirely new center.
@@
Two expert neuroradiologists annotated ten randomly sampled scans from the ISLES’22 training set [37]. When comparing their delineations against the ground truth masks, they achieved a median $\pm$ interquartile range Dice score of 0.92 $\pm$ 0.16 and a lesion-wise F1 score of 0.82 $\pm$ 0.30. The ensemble algorithm yielded, over the entire test set, a Dice score of 0.82 $\pm$ 0.12 and an F1 score of 0.86 $\pm$ 0.21. Figure 5 shows the median Dice MRI scan (or the closest to the median) for diverse stroke scenarios. The visual results of the predicted ischemic masks suggest that the ensemble predictions (green delineations) closely follow the manually segmented lesions (red delineations), exposing the algorithm’s capability to cope with diverse types of stroke patterns and configurations. The quantitative and qualitative results suggest robustness towards heterogeneous clinical and imaging scenarios.
@@
In a Turing-like test, nine experienced radiologists rated the stroke segmentation quality of the ISLES’22 test data. Radiologists received forty or forty-one randomized images with each image being delineated either by an expert or by the ensemble algorithm and were asked to rate the completeness and correctness of the lesion masks in a 1-to-6 (worst-to-best) scale. Boxplots of this experiment are shown in Fig. 6. Interestingly, the ensemble algorithm exhibits statistically significantly higher ratings than the experts (p-value = 0.02 when considering the segmentation completeness and p-value < 0.001 when considering the segmentation correctness, Wilcoxon signed-rank tests). The observation that experts find deep learning segmentations to be qualitatively superior to manually traced ones is unsurprising, given that similar findings were reported in prior research [40].
@@
We tested the ISLES’22 ensemble algorithm over the largest available acute and early sub-acute ischemic stroke cohort (N = 1686) retrieved from the Johns Hopkins Comprehensive Stroke Center [41]. The data were collected over ten years utilizing eleven magnetic field scanners operating at either 1.5T or 3T, sourced from four different vendors. Due to continual technological advancements over the last decade, the dataset comprises diverse MRI acquisition protocols and machine technologies, leading to strong variations in image quality (such as resolution, signal-to-noise ratio, and contrast-to-noise ratio). This makes this dataset uniquely suited to evaluate the generalizability of algorithms. Ischemic lesion masks were annotated by the dataset authors through manual delineation. Patient age was 62.5 $\pm$ 13.3 years. 907 patients (53.8%) were male. Reported race or ethnicity included 753 (44.7%) Black or African American, 490 (29.1%) White, and 40 (2.4%) Asian. 876 (52.0%) MRI scans were acquired after thrombolysis treatment with intravenous tissue-type plasminogen activator.